OpenStudy (anonymous):

The value of x = 22-18+98y-8(2) find the slope

OpenStudy (bibby):

Get it into y=mx+b form. Start by rewriting the equation and simplifying it into chunks. \[\large x = 22-18+98y-8(2)\] \[\large x = [22-18]+98y-[8*(2)]\]

OpenStudy (bibby):

so do the algebraic simplification within the brackets. 22-18 = ?? 8 * 2 = ??

OpenStudy (anonymous):

bibs tank yew su much u is suu kind man ;-; you r mi god nao ok okai

OpenStudy (bibby):

english pls megan

OpenStudy (bibby):

math pls megan*

OpenStudy (anonymous):

I forgot how to english math wut

OpenStudy (bibby):

Just simplify the equation\[\huge x = [22-18]+98y-[8*2]\]

OpenStudy (anonymous):

I really want to thank you and hug you right now, because man, you just saved my life ;-;

OpenStudy (bibby):

but we didn't even do the problem but I'll take the hug but still megan pls do the math